OpenStudy (anonymous):

Explain how you would find the exact value of 3 over the square root of 8.

OpenStudy (anonymous):

\[\frac{ 3 }{ \sqrt{8} }\]

OpenStudy (anonymous):

\[\frac{3}{\sqrt{8}}\]?

OpenStudy (anonymous):

if you want to write this is simplest radical form, multiply top and bottom by \(\sqrt{2}\)

OpenStudy (anonymous):

you get \[\frac{3}{\sqrt{8}}\times \frac{\sqrt{2}}{\sqrt{2}}=\frac{3\sqrt{2}}{\sqrt{16}}=\frac{3\sqrt{2}}{4}\]

OpenStudy (anonymous):

Where do you get sqrt(2) from?

OpenStudy (anonymous):

although... i have to add, there is nothing more EXACT about either form the second one is written in simplest radical form the first one is not

OpenStudy (anonymous):

i got it because it works if you multiply \(8\times 2\) you get a perfect square, namely \(16\)

OpenStudy (anonymous):

so \(\sqrt{8}\times \sqrt{2}=\sqrt{16}=4\) and you have no radical left in your denominator

OpenStudy (anonymous):

ahh okay. Understandable.

OpenStudy (anonymous):

now ask your math teacher the following question: why is \(\frac{3\sqrt{2}}{4}\) more EXACT than \(\frac{3}{\sqrt{8}}\) i would love to know the answer

OpenStudy (anonymous):

Good idea! :p

OpenStudy (anonymous):

what the question is really asking is "write \(\frac{3}{\sqrt{8}}\) in simplest radical form" which actually has a definition one condition is that there should be no radical in the denominator

OpenStudy (anonymous):

lol Satellite, "simplify" lol

OpenStudy (anonymous):

well "simplest radical form" actually means something

OpenStudy (anonymous):

true

OpenStudy (anonymous):

not like "simplify"

OpenStudy (anonymous):

i will bet $8 that @Mikeyy1992 math teacher is unclear as to what it means

OpenStudy (anonymous):

which is why the problem said "exact"

OpenStudy (anonymous):

I will inform exactly what she says (: